### v1

# Classical Mechanics from Energy Conservation or: Why not Momentum?

Abstract
It is demonstrated that energy conservation allows for a “heuristic” derivation of Newtonian mechanics, if the energy is presumed to be an additive function of position and velocity. It is shown that energy must be depicted as a function of position and momentum in order to allow for the correct relativistic equations. Accordingly it is argued that not only quantum theory but also special relativity is intrinsically Hamiltonian and requires a description by coordinates and momenta instead of coordinates and velocities.
Furthermore it is argued that the usual historical order of the “formulations” of mechanics, from Newtonian via Lagrangian to Hamiltonian mechanics, is illogical and misleading. Therefore it should be reversed.
I Introduction
The question whether and how it would be possibly to introduce mechanics from the concept of energy and energy conservation, received renewed interest in the recent years Hanc2004 ; LeGresley2019 . It has been suggested in this context to ignore traditional wisdom and to relate energy conservation to a “state function” of position and velocity Carlson2016 instead of the well-known methods of Hamiltonian mechanics. This attempt was criticized, but a clear exposition why it is generally required to express energy as a function of position and momentum (instead of position and velocity), has not been given Baumgarten2024 .
Here we aim to demonstrate in more detail why the use of energy conservation on the basis of coordinates and velocities allows to derive Newton’s laws of motion but fails in the “relativistic” case. We hope to provide sufficient evidence for the claim that relativistic mechanics – when derived from energy conservation – is intrinsically Hamiltonian.
II A Quantitative Concept of Work
There is the tale of Galileo climbing up the tower of Pisa for an experimental test of free fall GUNTHER1935 . The story tells that the experiments provided evidence that free fall is independent of the mass (and other properties) of the falling objects. Let’s assume (for the sake of the argument) that Galileo had also a (perfect) rubber ball at hand and that he observed that this rubber ball was not just accelerated towards the ground, but that it was also reflected and returned to (almost) the same height. This way Galileo might have realized that the work he invested into the ball’s height was converted into motional energy during the fall, into elastic deformation on impact, back into motional energy when reflected and eventually back into positional work when returned to the height of the tower. He was able to understand that a feather and some iron ball would fall equally rapid in vacuum, and he might have been able to understand that work is conserved in his experiment provided that friction can be neglected.
If Galileo would have had the appropriate knowledge of (vector-) calculus (which was later invented by Newton and Leibniz), he could have derived Newton’s laws from the simple idea that the sum of the work stored in position $V(x)$ and the work stored in motion must be constant, provided we can neglect friction. Since motion is described by the velocity $v={dx\over dt}$, the corresponding work stored in motion $T(v)$ should depend primarily on the velocity – or on some other “quantity of motion”.
III Energy as a Function of Position and Velocity
Hence Galileo could have started from the observation that stored work appears in (at least) two forms, namely in the form of positional work $V(x)$ and in the form of motional work $T(v)$. These two forms can be transformed into each other but the total work, the sum of the two contributions, is (approximately) conserved. We may then express this condition of the conservation of work (or energy $E$, respectively) by
$\begin{split}E(x,v)&=T(v)+V(x)\\ {dE\over dt}&={\partial E\over\partial v}\,{dv\over dt}+{\partial E\over\partial x}\,{dx\over dt}=0\\ &={dT\over dv}\,{dv\over dt}+{dV\over dx}\,v=0\,.\end{split}$ (1)
Division of the last line by the velocity $v$ and acceleration ${dv\over dt}$ gives:
${1\over v}\,{dT\over dv}=-{{dV\over dx}\over{dv\over dt}}\,.$ (2)
The left side of Eq. 2 is exclusively a function of the velocity (and possibly of properties of the moving object), but the velocity does not appear on the right side. Hence, if the equation holds for all velocities, then both sides must be equal to some constant factor; let’s call it $m$:
$\begin{split}m&={1\over v}\,{dT\over dv}\\ m&=-{{dV\over dx}\over{dv\over dt}}\\ \end{split}$ (3)
The first equation can be directly integrated:
$\begin{split}{dT\over dv}&=m\,v\\ T&=\frac{1}{2}\,m\,v^{2}+C\\ \end{split}$ (4)
where the integration constant $C$ can be set to zero, since only the derivative of $T(v)$ is used in the equations of motion.
The (classical) work stored in motion is then uniquely determined. Furthermore, if one now introduces the acceleration $a={dv\over dt}$ and the defines the “force” by $F=-{dV\over dx}$ and inserts this into the second equation, then Newton’s second law appears:
$F=m\,a$ (5)
This Ansatz directly and inevitably leads to Newtonian mechanics with $F=m\,a$ and $T={m\over 2}\,v^{2}$. This means that Newton’s “laws of motion” can be obtained, at least in large part, from a concept of energy conservation, when expressed as a function of position and velocity.
Unfortunately this implies that it can not possibly be used to obtain the correct (“relativistic”) equations of dynamics.
IV Energy as a Function of Position and Momentum
In the previous case we simply presumed that position and velocity can be treated as fully independent variables. This seemingly innocent assumption provided the possibility to derive a definite expression for the kinetic energy $T={m\over 2}\,v^{2}$, independent of the functional form of the positional energy (i.e. potential).
If we are not sure whether this is correct in all cases, then one can reduce the implications of this assumption and assume merely that the work of motion must depend on some quantity of motion $p$. It is only presumed that this quantity has a monotonic functional relationship to the velocity.
Then one may proceed as before: The energy $E$ is the sum of two contributions, namely the work stored in position $V(x)$ and the work stored in motion $T(p)$:
$\begin{split}E(x,p)&=T(p)+V(x)\\ {dE\over dt}&={\partial E\over\partial p}\,{dp\over dt}+{\partial E\over\partial x}\,{dx\over dt}=0\\ {dE\over dt}&={dT\over dp}\,{dp\over dt}+{\partial V\over\partial x}\,v=0\\ \end{split}$ (6)
again with $v={dx\over dt}$. Division by $v$ and ${dp\over dt}$ gives
${1\over v}\,{dT\over dp}=-{{\partial V\over\partial x}\over{dp\over dt}}\,.$ (7)
The left side depends (again) on $v$ and $p(v)$ alone. Hence it must be equal to some constant, which can now be freely chosen to be equal to one, since the scale and unit of $p$ are not yet fixed and can be chosen accordingly.
This means that we are free to call the term $-{\partial V\over\partial x}$ force $F$ and hence this term must be equal to ${dp\over dt}$. Then we obtain the general (Hamiltonian) “definition” of velocity:
$v={dT\over dp}\,.$ (8)
This time we do not obtain a definite expression for $T(v)$ or $T(p)$. And this implies that it covers a wider range of possibilities. Hence this way we obtain a relationship that is not limited to Newtonian mechanics. The mere fact that a more general case is possible already implies that energy should be regarded as a function of coordinate and momentum unless additional information is available. But it also means that we don’t have the complete solution, since the last equation cannot be directly integrated as long as we don’t know how $p$ and $v$ are related.
It is rather easy to argue that $T(p)$ (or $T(v)$, respectively) must be an even function of it’s argument, namely because the sign of the “momentum” $p$ may not alter $T(p)$, so that the lowest order term that is relevant to the equations of motion must be the quadratic term. We can then write the Taylor series of this function as follows:
$T(p)=\frac{a_{2}}{2!}\,p^{2}+\frac{a_{4}}{4!}\,p^{4}+\dots$ (9)
Combined with Eq. 10 this yields:
$v={dT\over dp}=a_{2}\,p+\frac{a_{4}}{6}\,p^{3}+\dots$ (10)
Hence the first non-trivial term, that can be used in good approximation, is
$v\approx a_{2}\,p$ (11)
so that, as before, only the determination of the constant $a_{2}=m^{-1}$ is left as the remaining step to arrive at Newton’s “laws”. The obvious advantage of this latter approach is that it uncovers a hidden (and wrong) assumption of Newtonian mechanics and a possible cure, namely the use of momentum instead of velocity as a primary variable.
V Lagrangian Mechanics
Another advantage of the use of the “quantity of motion”, namely the momentum $p$, is the possibility to derive Lagrangian mechanics via the Legendre transform. Since, if we use the variables $v$ and $p$ and make no apriori assumption about their relationship, then the Legendre transform allows to derive the Lagrangian ${\cal L}(x,v)$ from the “Hamiltonian” $E={\cal H}(x,p)$:
${\cal L}(x,v)=v\,p-{\cal H}(x,p)\,.$ (12)
The momentum $p$ can then obviously be defined by
$p={\partial{\cal L}\over\partial v}\,.$ (13)
The second partial derivative is then
${\partial{\cal L}\over\partial x}=-{\partial{\cal H}\over\partial x}=-{dV\over\partial x}$ (14)
and the equation of motion is therefore:
${dp\over dt}={d\over dt}\left({\partial{\cal L}\over\partial v}\right)={\partial{\cal L}\over\partial x}\,,$ (15)
or, respectively:
${d\over dt}\left({\partial{\cal L}\over\partial v}\right)-{\partial{\cal L}\over\partial x}=0$ (16)
Obviously this is a reversal of the usual textbook order insofar as we derive Lagrangian from Hamiltonian dynamics and not vice versa. This order uncovers that the so-called “principle of least action” (PLA) is not needed to derive the Lagrangian equation(s) of motion.
It is tradition to introduce classical mechanics in historical order; it invariably starts with Newton’s “laws”, proceeds with Lagrangian mechanics and finally arrives via the Legendre transform at the Hamiltonian “formulation”. These three “formulations” of classical mechanics are often said to be equivalent Davis1986 ; Kelley2017 - which is false as just demonstrated.
Many textbook authors, also dinstinguished and renowned physicists like Feynman and co-authors FLS , introduce Lagrangian mechanics from the PLA, which is depicted as “absolutely fascinating”, as a mystifying and “deep” principle of nature, echoing some of the theologically and/or philosophically motivated scientists of the 18th and 19th centuries, who thought they had found a deep and significant principle of nature. We quote (part of) the conclusion of Yourgrau and Mandelstam Yourgrau :
Obsessed with the urge to achieve ”cognition of real nature, ” i.e., to gain an ”explanation” of material phenomena, physicists have assigned to a particle the knowledge which enables it to take the most convenient path of all those which would direct it to its destination. The Aristotelian final cause is dragged into the subject, and the particle is presented to us as what Poincare calls ”a living and free entity.” Aristotle’s dictum that nature does nothing in vain may have been at the foot of Maupertuis’ teleological principle, which initiated the study of variational principles in nature, but to the twentieth-century physicist the notion of such a mysterious purposive agency would only be a regress into sheer obscurantism. History knows of many instances where a science had its beginnings in a mythical and magical symbolism which gradually shed its mystical features as our rational knowledge increased and our conceptions became clearer- witness the transition from astrology to astronomy and from alchemy to chemistry.
But the issue is not just of philosophical nature. The Lagrangian is usually introduced as ${\cal L}=T-V$. This however is wrong in (relativistic) mechanics proper, and hence the traditional lore invariably has to move the goal posts when one wishes to “derive” the relativistic equations of motion from a Lagrangian.
It should be noted, however, that the goal posts remain untouched when classical mechanics is taught not in historical but in logical order, namely if the Lagrangian is derived from the Hamiltonian.
VI “Relativistic” Mechanics
From the perspective of pre-relativistic mechanics, the freedom to use other functions for $T(p)$ than $T=p^{2}/(2\,m)$ seems to be of no use. In the times of Galileo and Kepler, experiments would have quickly lead to the conviction that terms beyond second order in $T(p)$ have no physical relevance. Today we know that this is wrong: The fact that $T(p)$ can not be completely determined from the condition of energy conservation is indeed physically significant. It allows to reconcile classical mechanics with relativity.
The proper “relativistic” energy-momentum-relation, also called dispersion relation for a free particle is well-known:
$E=\sqrt{m^{2}\,c^{4}+c^{2}\,{\bf p}^{2}}\,.$ (17)
This equation was not but could have been found simply by measurements: Modern accelerators cover many orders of magnitude of the argument ${\bf p}/(m\,c)$ so that we could obtain Eq. 17 entirely from a fit to the data 111 Special relativity was necessary to unify mechanics and electrodynamics into a single coherent framework. Insofar relativistic dynamics was derived from the invariance of electromagnetism under Lorentz transformations. However there is also a theoretical possibility to “derive” the Dirac-Clifford algebra and in consequence the form of Eq. 17 and the Lorentz transformations from an algebraic analysis of a pure and abstract Hamiltonian Ansatz qed_paper ; osc_paper , i.e. from a “pure” conservation law. A discussion of this approach would exceed the scope of this paper. .
Of course, in order to build modern high-energy accelerators, one has to make use of Eq. 17 so that this argument is somewhat hypothetical  [^2] Nonetheless the fact that the use of this equation allows for the successful construction and design of particle accelerators provides hard evidence for it’s validity.. The velocity is then given not as a primary but as a quantity derived from Eq. 10 (one of Hamilton’s equations of motion):
$v=\nabla_{p}E({\bf p})={c^{2}\,{\bf p}\over\sqrt{m^{2}\,c^{4}+c^{2}\,{\bf p}^{2}}}\,.$ (18)
With some basic algebra, this leads to
$\begin{split}E&=m\,c^{2}\,\gamma\\ \vec{p}&=m\,c\,\gamma\,\vec{\beta}\\ \end{split}$ (19)
with the usual $\vec{\beta}=\vec{v}/c$ and $\gamma=(1-\beta^{2})^{-1/2}$.
There is another formal reason why relativistic dynamics is intrisically Hamiltonian. As a matter of fact, velocities are (in the correct relativistic theory) not additive. The co-called relativistic addition formula is a paradoxical misnomer, since if velocities were additive it would not need a formula to add them. But the momentum is additive in any version of mechanics, relativistic or not. Hence the momentum vector is really a member of a mathematical vector space, while velocity is not.
VII Conclusions
The first principle is that you must not fool yourself ­­ and you are the easiest person to fool … you should not fool the layman when you’re talking as a scientist.
– Richard Feynman Joking
We hold that the term “relativistic mechanics” is a misnomer: Classical mechanics proper is always “relativistic”, because this is the general case. Only if one restricts the considerations to the Newtonian approximation one might legitimately use the more specific label “non-relativistic”. This is not only the perspective from accelerator physics, but it is also the natural perspective if one aims for a straight deductive account of mechanics.
Unfortunately the usual textbook introduction to mechanics is not deductive, as it should be, but an excercise in epigonism: Inevitably common textbook mechanics begins with Newton’s “laws”, “derives” Langrangian mechanics from it and finally “derives” Hamiltonian from Lagrangian mechanics. This is the standard approach and only a minority of textbooks and lecture notes do not simply copy-paste this approach.
Though this account mimics the historical order, it is rarely sufficiently emphasized that Lagrangian is more general than Newtonian mechanics and Hamiltonian is more general than Lagrangian mechanics. Unless one wishes to teach history and not physics, a deductive approach should always start from the most general laws or facts and then derive the more specific from the more general laws. Of course, this has been suggested before, for instance by Hanc and Taylor Hanc2004 .
If a Hamiltonian approach is available, the Lagrangian approach can be obtained from the former via the well-known Legendre transformation. If one starts from the conservation of energy (using momentum and not velocity!), i.e. from Hamiltonian mechanics, then one always obtains the “correct” Langrangian, while if one mimics the historical order, then one is obliged to explain why the “classical” definition of the Lagrangian, namely ${\cal L}=T-V$, does not hold in the “relativistic case”. Again: This problem does not arise in the logically correct order which starts with Hamiltonian mechanics.
Furthermore, to start with the Hamiltonian helps to preserve some scientific sobriety, since in this case Lagrangian mechanics is “derived” and explicitely not fundamental. Thus the PLA would not be depicted any more as a mystifying “deep” principle of nature but rather as a theorem of mathematical physics that can be (and has to be) derived from energy conservation.
It is not at all surprizing that a conservation “law” is accompanied by an extremum: A curve with constant bending radius is a circle; and it happens to be the case that a circle is the shape with the least circumference for a given area. This is simply a fact of geometry and not a “deep” and mystifying principle of nature. Furthermore, as Yourgrau and Mandelstam already noted many years ago Yourgrau :
Untenable as the arguments in support of teleological ideas underlying variational principles have proved, it still remains for us to inquire into the so-called all-embracing nature of these laws. Now it cannot be contested that variational principles do enable us to condense several mathematical equations describing physical phenomena into one compact theorem; furthermore we behold here one type of principle which governs the laws of a large part of physics. This fact is, however, by no means extraordinary, because these branches of physics are all characterized by differential equations, and it is not at all an exceptional property of a set of differential equations to be expressible as a variational principle.
Richard Feynman warned us not to fool ourselves. As he claimed, it is easy to fool yourself and then, as a professor, to fool students and/or laypeople. But in their famous and celebrated lecture series, Feynman, Leighton and Sands (FLS) spent a considerable number of pages and figures to introduce and explain the PLA and the Lagrangian, defined by ${\cal L}=T-V$ FLS . However, after discussing the meaning of ${\cal L}=T-V$, all textbooks have to move the goal posts for the “relativistic” case in order to explain why the “relativistic” Lagrangian is not the difference between kinetic and potential energy. It is an easy exercise to verify that the relativistic Hamitonian directly provides the correct relativistic Lagrangian. Even (and especially) a layman might thus get the impression that physicists actually fool themselves regarding the PLA 333 See for instance the rather recent homages to the PLA Hone2004 ; Coopersmith2017 . In Ref. Gray2007 we read that “Several authors have simplified and elaborated the action principle and recommended that it be introduced earlier into the physics curriculum”. We argue to the contrary that it should, if at all, be introduced as a consequence of energy conservation, that is: never before Hamiltonian mechanics. : The PLA is possibly intriguing and fascinating, but it has no deep physical significance whatsoever.
Firstly the “principle” refers to the least action though the action is not always the least Yourgrau ; Gray2007 . It has been suggested to call it instead the “principle of stationary action” (PSA) Gray2007 . But this again moves the goal posts, since if the action is not the least but only “stationary”, then the original claim that the PLA provides a proof for an “optimal world” can hardly be defended. Secondly, the principle requires more explanations than it can possibly provide, because even if one finds that some action functional is stationary, this is of little practical value unless it is possible to derive the action functional from fundamental principles. And this is exactly what we suggest here: we have a principle at hand that - to our best knowledge - holds without any exception, namely the principle of energy conservation. Not only that this latter “principle” refers to a measurable physical quantity and is easy to explain and understand, one can furthermore derive the (correct) Lagrangian from it. It can then be demonstrated that the Lagrangian provides an action functional that is stationary. This can be done without moving the goal posts and without mystification.
Energy conservation might seem “trivial” when compared to the “intriguing” PLA, but we should not fool ourselves: A quote attributed to E. Rutherford claims that “all of physics is either impossible or trivial. It is impossible until you understand it, and then it becomes trivial”. As it seems the PLA never became trivial and this might be an indication that it was never “understood”. We suspect that it was never understood because there is nothing about this “principle” that could be understood. This is not meant to downplay Lagrangian mechanics as a mathematical method, which can be derived in a few steps from energy conservation. But to insist that the PLA or PSA or any other “optimality condition” provides insight into a deep “principle of nature” is just wishful thinking and an efficient method to fool yourself.
Energy conservation is not a property of an “optimal” but rather a property of a real world. It’s trivialized message is that there is no free lunch: If you want work to be done, you have to pay with energy for it. In the view of the author it is a necessary truth in any real world. Only in our dreams we may (and probably do) forget about it.